You expect KT industries​ (KTI) will have earnings per share of $ 4 this year and expect that they will pay out $ 1.75 of these earnings to shareholders in the form of a dividend. ​ KTI's return on new investments is 14​% and their equity cost of capital is 10​%. The value of a share of​ KTI's stock is closest​ to:

A.$32.94

B.$98.82

C.$82.35

D.$ 49.41
Answer #1

C. $82.35

g = retention rate × return on new investment = (4.00 - 1.75) / 4.00 × 0.14 = 0.07875 or 7.875%

P0 = Div1 / (rE - g) = 1.75 / (0.10 - 0.07875) = $82.35